OpenStudy (anonymous):

Help Please!!!!!!!!!! Solve these equations using the addition property. 2x + y = -2 x + y = 5 The lines intersect at ( __, __ )

OpenStudy (anonymous):

2x + y = -2 x + y = 5 _____________ 2x + y = -2 -2x - 2y = -10 ______________ -y=-12 y=12 -----> x + 12 = 5 ---> x=-7 (-7, 12)